# RECAST: Strengthening LLMs’ Complex Instruction Following with Constraint-Verifiable Data

### v1

@@ 5.1 Setups
We constructed RECAST-Test, a hierarchical evaluation benchmark based on RECAST-30K. This benchmark includes four different difficulty levels, categorized by increasing constraint complexity. Each level comprises instructions with progressively more constraints, enabling fine-grained performance evaluation across tasks of varying complexity. The detailed construction method and statistical characteristics of each difficulty level are provided in Appendix A.2. To assess the generalization capabilities of RECAST beyond our constructed benchmark, we further evaluated it on FollowBench [21], a multi-level benchmark that includes five detailed constraint categories: Content, Situation, Style, Format, and Example.
@@ 5.2.1 Evaluation on RECAST-Test
Table 2: Results of FollowBench across different levels. The best result in each column is highlighted in bold, while the second-highest result is indicated with underline formatting.
@@
Level 1 Level 2 Level 3 Level 4 Level 5 Level 1 Level 2 Level 3 Level 4 Level 5
  Conifer $64.41$ $51.06$ $44.31$ $35.77$ $27.92$ $44.69$ $78.24$ $62.91$ $46.50$ $43.95$ $29.76$ $52.27$
Crab $65.98$ $52.72$ $38.44$ $31.40$ $20.96$ $41.90$ $73.01$ $53.82$ $39.11$ $33.15$ $23.41$ $44.50$
I-SHEEP $51.55$ $41.15$ $28.16$ $17.99$ $8.55$ $29.48$ $61.32$ $41.93$ $31.98$ $19.19$ $22.50$ $35.38$
MUFFIN $25.59$ $14.72$ $12.74$ $17.07$ $13.78$ $16.78$ $62.40$ $58.00$ $42.72$ $26.80$ $20.85$ $42.15$
ShareGPT $69.93$ $50.45$ $39.91$ $39.26$ $27.49$ $45.41$ $75.66$ $61.53$ $42.79$ $37.80$ $27.54$ $49.06$
Evol-Instruct $56.96$ $46.77$ $34.01$ $27.47$ $17.81$ $36.60$ $71.92$ $60.04$ $42.66$ $30.51$ $26.58$ $46.34$
Suri $5.66$ $2.67$ $5.15$ $3.42$ $3.20$ $4.02$ $7.77$ $4.87$ $5.62$ $4.99$ $4.68$ $5.59$
Tülu 3 Persona IF $75.53$ $59.98$ $57.85$ $40.09$ $43.75$ $55.44$ $82.81$ $71.02$ $51.20$ $44.63$ $35.02$ $56.94$
RECAST-30K-SFT $78.26$ $67.16$ $51.15$ $49.02$ $39.91$ $57.10$ $79.20$ $69.49$ $55.87$ $48.72$ $44.71$ $59.60$
RECAST-30K-RLVC $77.86$ $70.12$ $56.65$ $54.32$ $49.87$ $61.76$ $78.21$ $71.37$ $57.25$ $56.12$ $53.21$ $63.23$
5.2.2 Evaluation on FollowBench
Table 2 presents performance results on FollowBench, demonstrating the strong generalization capabilities of our approach. RECAST-30K-SFT shows substantial improvements over all baseline methods, achieving average performance gains of $1.66$% and $2.66$% over the strongest baseline (Tülu 3 Persona IF) on Llama3.1-8B and QWEN2.5-7B, respectively. RECAST-30K-RLVC further enhances these results, with particularly significant gains at higher difficulty levels (Levels 4-5), where constraint complexity increases. The consistent performance advantage across this out-of-domain benchmark confirms that models fine-tuned with RECAST effectively learn generalizable instruction-following capabilities rather than merely overfitting to specific constraint patterns in the training data.
@@ C.1 Baseline
C.2 Details of Data Generation

### v2

@@ 5.1 Setups
We constructed RECAST-Test, a hierarchical evaluation benchmark based on RECAST-30K. This benchmark includes four different difficulty levels, categorized by increasing constraint complexity. Each level comprises instructions with progressively more constraints, enabling fine-grained performance evaluation across tasks of varying complexity. The detailed construction method and statistical characteristics of each difficulty level are provided in Appendix A.2. To assess the generalization and general capabilities of RECAST beyond our constructed benchmark, we further evaluated it on six benchmarks, the details of these benchmarks can be found in Appendix C.2.
@@ 5.2.1 Evaluation on RECAST-Test
Table 2: Results of multiple benchmarks across different models. IFEVAL uses Prompt-Level Loose Accuracy (Pr. (L)), measuring the proportion of prompts where all instructions are followed using a loose criterion; FollowBench uses Hard Satisfaction Rate (HSR), evaluating complete constraint satisfaction across multi-level instructions; BBH, GPQA, MUSR, and MMLU-PRO use standard Accuracy (Acc), measuring the percentage of correct responses. The best result in each column is highlighted in bold, while the second-highest result is indicated with underline formatting.
@@
IFEVAL Followbench BBH GPQA MUSR MMLU-PRO IFEVAL Followbench BBH GPQA MUSR MMLU-PRO
Pr. (L) HSR Acc Acc Acc Acc Pr. (L) HSR Acc Acc Acc Acc
  Conifer $43.62$ $44.69$ $35.49$ $27.61$ $37.64$ $12.89$ $33.66$ $44.73$ $52.27$ $49.80$ $31.22$ $45.54$ $40.94$ $44.08$
Crab $42.14$ $41.90$ $46.85$ $32.48$ $40.26$ $32.18$ $39.30$ $44.55$ $44.50$ $49.16$ $31.89$ $48.74$ $41.72$ $43.43$
I-SHEEP $32.90$ $29.48$ $43.63$ $28.33$ $36.30$ $29.95$ $33.43$ $36.60$ $35.38$ $45.44$ $29.95$ $43.18$ $43.09$ $38.94$
MUFFIN $24.95$ $16.78$ $28.77$ $26.64$ $33.69$ $11.66$ $23.75$ $36.78$ $42.15$ $29.42$ $29.07$ $37.93$ $15.92$ $31.88$
ShareGPT $46.40$ $45.41$ $45.96$ $30.02$ $40.16$ $30.09$ $39.67$ $47.13$ $49.06$ $49.00$ $31.20$ $44.95$ $40.13$ $43.58$
Evol-Instruct $42.70$ $36.60$ $44.31$ $28.41$ $41.75$ $28.03$ $36.97$ $45.29$ $46.34$ $51.33$ $32.55$ $47.41$ $43.18$ $44.35$
Suri $18.30$ $4.02$ $29.50$ $26.09$ $33.40$ $11.81$ $20.52$ $12.75$ $5.59$ $29.23$ $26.60$ $41.79$ $11.69$ $21.28$
Tülu 3 Persona IF $73.94$ $55.44$ $30.10$ $28.64$ $38.68$ $12.50$ $39.88$ $73.94$ $56.94$ $44.75$ $31.06$ $44.13$ $41.32$ $48.69$
RECAST-30K-SFT $76.34$ $57.10$ $29.85$ $26.77$ $37.39$ $11.72$ $39.86$ $73.38$ $59.60$ $38.87$ $30.02$ $39.91$ $37.87$ $46.61$
RECAST-30K-RLVC $77.39$ $61.76$ $30.03$ $30.81$ $38.47$ $12.12$ $41.76$ $74.01$ $63.23$ $34.69$ $29.84$ $46.94$ $37.07$ $47.63$
5.2.2 General Capability Evaluation
Table 2 evaluates RECAST’s impact on general model capabilities across six diverse benchmarks. Despite training on only 30K examples focused on constraint-following, RECAST-30K-RLVC demonstrates strong performance across all methods, achieving scores of $41.76$ on Llama3.1-8B and $47.63$ on QWEN2.5-7B. The models demonstrate strong performance not only on instruction-following tasks such as IFEVAL with $77.39$ and Followbench with $61.76$, but also maintain competitive results on reasoning benchmarks, including BBH and MuSR, as well as knowledge evaluation tasks like GPQA and MMLU-PRO. This balanced improvement across fundamentally different task types confirms that RECAST’s constraint-verifiable training approach enhances models’ general capabilities without sacrificing performance on traditional benchmarks, validating its effectiveness as a comprehensive training methodology.
@@ C.1 Baseline
C.2 General Capability Evaluation Benchmark
We employ six diverse benchmarks spanning different aspects of language understanding and reasoning:
  • FollowBench [21] comprehensively evaluates instruction-following capability through five types of fine-grained constraints (Content, Situation, Style, Format, and Example) using a multi-level mechanism that incrementally adds constraints to initial instructions. This benchmark systematically measures models’ ability to satisfy multiple constraints simultaneously, with difficulty levels ranging from simple single-constraint instructions to complex multi-constraint scenarios. The evaluation employs metrics including Hard Satisfaction Rate (HSR) for complete constraint satisfaction and Soft Satisfaction Rate (SSR) for partial compliance.
  • IFEval [45] assesses models’ ability to follow explicit formatting instructions through strict, verifiable metrics rather than subjective content evaluation. The benchmark focuses on precise adherence to specific requirements such as keyword inclusion, format specifications, and structural constraints. This allows for rigorous, code-based evaluation that eliminates ambiguity in measuring instruction compliance.
  • BBH [46] comprises 23 carefully selected challenging tasks from the BigBench dataset that test diverse cognitive capabilities including multistep arithmetic, algorithmic reasoning, language understanding, and world knowledge. These tasks are chosen for their high difficulty, objective metrics, and strong correlation with human preferences. The benchmark provides insights into models’ reasoning abilities across domains ranging from boolean expression evaluation to sarcasm detection.
  • GPQA [47] features expert-crafted questions from PhD-level domain specialists in biology, physics, and chemistry designed to be challenging for non-experts yet accessible to specialists. The dataset undergoes rigorous validation to ensure both factual accuracy and appropriate difficulty levels. Access restrictions and gating mechanisms protect against data contamination, maintaining the benchmark’s integrity for evaluating advanced scientific knowledge.
  • MuSR [48] tests models’ ability to integrate reasoning with long-context understanding through algorithmically generated complex problems averaging 1,000 words in length. The benchmark includes murder mysteries, object placement puzzles, and team allocation optimization tasks that require maintaining coherent reasoning across extended contexts. Few models achieve above-random performance, making it an effective discriminator of advanced reasoning capabilities.
  • MMLU-PRO [49] represents an enhanced version of the original MMLU benchmark, addressing previous limitations through increased difficulty and improved quality control. The benchmark presents 10 answer choices instead of 4, requires deeper reasoning, and underwent expert review to eliminate noisy or unanswerable questions. This refinement results in a more challenging and reliable assessment of multitask language understanding across diverse knowledge domains.
C.3 Details of Data Generation